OpenStudy (truwhovian11):

first person to answer my question gets a fan medal and a testimonial A store charges $8.95 for an 8 ounce jar of maple syrup. The price for a 10 ounce jar of the same brand is $12.00. Which jar is the better buy? What is the difference in the unit prices? If necessary, round to the nearest cent. A) 8 ounce jar; $0.08 B) 8 ounce jar; $0.15 C) 10 ounce jar; $.12 D) 10 ounce jar; $0.08

OpenStudy (mathstudent55):

Divide the price by the number of ounces. That gives you the unit price.

OpenStudy (truwhovian11):

@-SpencerBraz- @ace-n-it @Awolflover1 @baru @drawwithapurpose @Hauntedwoodsgal @helpkay @logan13 @niyex @ojal26 @Roxy.girl @welshfella @YoungStudier

OpenStudy (mjdennis):

Do you know how to find out how much each item costs for 1 ounce?

OpenStudy (truwhovian11):

no

OpenStudy (helpkay):

Like @mathstudent55 said divide the price with the amount of juice given

OpenStudy (truwhovian11):

is it C

OpenStudy (mathstudent55):

A better buy is an item that costs less per unit.

OpenStudy (mathstudent55):

$8.95 for an 8 ounce: $8.95/(8 oz) = $1.12/oz 10 ounce jar of the same brand is $12.00: $12/(10 oz) = $1.20/oz The 8-oz jar is the better deal since $1.12/oz is less expensive than $1.20/oz. The difference is: $1.20/oz - $1.12/oz = $0.08/oz

OpenStudy (mrnood):

@mathstudent55 so - having given a good method for the poster to follow (and nothing but a guess in return) WHY do you post the complete answer? It is not allowed, and it doesn't help the poster to learn

OpenStudy (xmissalycatx):